OpenStudy (anonymous):

Find all the zeroes of the equation. –3x4+ 27x2 + 1200 = 0

OpenStudy (anonymous):

-3(x^4 - 9x^2 - 400) = 0

OpenStudy (anonymous):

So it is because the GCF is -3??

OpenStudy (anonymous):

yea....

OpenStudy (anonymous):

I know the real roots are -5 and 5. Don't I need that to find out??

OpenStudy (anonymous):

so they already gave you these roots? u shouldve put put them along with the question

OpenStudy (anonymous):

use synthetic division then!

OpenStudy (anonymous):

use that synthetic division to find the other roots

OpenStudy (anonymous):

No, I figured out the roots with a graphing calculator and my answers were off after doing synthetic division.

OpenStudy (anonymous):

ok!!

OpenStudy (anonymous):

Alright. I'll just re-do the problem and thanks for the help( :

OpenStudy (anonymous):

om wait! lets work it out tpgether!

OpenStudy (anonymous):

Alright.

OpenStudy (anonymous):

do you know how to find rational zeros?

OpenStudy (anonymous):

So I need to figure out the zeros of the equation –3x4+ 27x2 + 1200 = 0. In my book it tells me that I need to find the real roots which are -5 and 5 and then use the synthetic division.

OpenStudy (anonymous):

I'm sure I would if I saw how to but not really.

OpenStudy (anonymous):

oh ok! -3(x^4 - 9x^2 - 400) = 0 -3(x^2 - 25)(x^2 + 16) = 0

OpenStudy (anonymous):

Set each factor equal to zero

OpenStudy (anonymous):

x^2 - 25 = 0 x = 5, -5

OpenStudy (anonymous):

Ohh! This is where I get confused...what do I need to do with -3?

OpenStudy (anonymous):

x^2 + 16 = 0 x = 4i, -4i

OpenStudy (anonymous):

Just leave that out!!!

OpenStudy (anonymous):

Because when you set =3 equal to zero: -3 = 0 they are NOT equal to zero!!!!!

OpenStudy (anonymous):

But when you simplify, you should write it out as: -3(x^2 + 16)(x-5)(x+5) = 0

OpenStudy (anonymous):

Nope! They aren't equal to 0. And I get it now!! Thanks for your help( :

OpenStudy (anonymous):

Its ok!! Thank you for the medal!!!!!!! :)

OpenStudy (anonymous):

No problem( :

OpenStudy (anonymous):

:)